OpenStudy (anonymous):

Use the integral test to determine if the series is convergent or divergent 9/2x-1

OpenStudy (anonymous):

\[\int\limits_{t}^{1}\frac{ 9 }{ 2x-1 }\]

OpenStudy (anonymous):

woops, t and 1 are switched

OpenStudy (anonymous):

I got the integral as being equal to 9(ln(2t-1))

OpenStudy (loser66):

yup

OpenStudy (anonymous):

My homework software is saying that's wrong :(

OpenStudy (loser66):

I think you should post the original problem, since you say "the series converge or not" I assume that you have something else

OpenStudy (anonymous):

I tried even including the 9(ln1) just incase

OpenStudy (loser66):

the integral give out the number, value of something, not give you the conclusion of converge or diverge. If it is the series or sequence, may be..... who know? may be I can understand what happen

OpenStudy (anonymous):

\[\sum_{0}^{\infty} \frac{ 9 }{ 2n-1 }\]

OpenStudy (loser66):

If I don't make mistake, the integral test says "If f is a continuous , positive, decreasing function on [1,infinitive, and \(a_n\) = f(n) for all n, then \(\sum_ {n=1}^{\infty} a_n and \int_1^{\infty} f(x)dx \) both converge or diverge remember 1 to infinitive, not 1 to t

OpenStudy (anonymous):

We were taught to evaluate the limit from 1 to t as t approached infiniti

OpenStudy (loser66):

nope, you confuse with the concept of take limit of the inpropriate integral.

OpenStudy (anonymous):

But that's even what our assignment says to do

OpenStudy (loser66):